OpenStudy (anonymous):

Earl Miller deposited $25,000 at Y Bank at an interest rate of 12% compounded quarterly. (Use the tables in the handbook.) The effective rate (APY) is

OpenStudy (jdoe0001):

\(\bf APY=\left(1+\frac{r}{n}\right)^n \\ \quad \\ r=rate\to 12\%\to \frac{12}{100}\to 0.012\\ n=\textit{compounding cycle, quarterly, means 4 times per year}\to 4\)